10. Problems and Applications Q4 Suppose that there are 10 million workers in Canada and that each of these workers can produce either 2 cars or 30 bushels of wheat in a The opportunity cost of producing a car in Canada isbushels of wheat, and the opportunity cost of producing a bushel of wheat in Canada is cars. Use the blue line (circle symbol) to draw Canadas production possibilities frontier (PPF) on the following graph. Then use the black point (plus symbol) to indicate the consumption bundle Canada can achieve without trade if it chooses to consume 10 million cars. 18 PPF Without Trade With Trade 0 50 100 50 200 250 300 30 400
20 T PPF 12 Without Trade 10T With Trade 0 50 100 150 200 250 300 350 400 Wheat (millions of bushels) Now suppose that the United States offers to buy 10 million cars from Canada in exchange for 20 bushels of wheat per car. On the previous graph, use the grey point (star symbol) to indicate the consumption bundle Canada can achieve with trade if it continues to consume 10 million cars Canada accept the deal the United States proposes. Save & Continue
Answer #1

Consider the given problem here there are “10 million” workers each can produce either “2 cars” or “30 bushels of wheat”, => the labor requirement for both the goods are given by “1/2 = 0.5” and “1/30”. So, the PPF is given by.

=> C/2 + W/30 = 10 (in million), where “C=cars” and “W=bushels of wheat”.

The opportunity cost of producing a car in Canada is “(1/2)/(1/30) = 15” bushels of wheat and the opportunity cost of producing a bushels of wheat in Canada is “1/15” cars. Consider the following fig.

Car 20 million E1 10 million 75 million 100 million 300 million Wheat in bushels

Now, Canada chooses to consume “10 million cars”, => “C=10”, => C/2 + W/15 = 10, => 10/2 + W/15 = 10.

=> 5 + W/15 = 10, => W = 5*15 = 75 bushels of wheat.

Now, suppose that the US offer to buy 10 million cars from Canada in exchange for “20 bushels” of wheat per car, => US will export “10 million cars” in exchange for “10*20 = 200 million bushels of wheat”. So, here if Canada specialize to “wheat”, => will be able to produce total “300 million bushels of wheat” and will export “200 million bushel, => will remain “100 million bushels”. So, here Canada will consume “10 million cars” and “100 million bushels of wheat”.

So, in the above fig we can after trade Canada continue to consume “10 million cars”. So, the Canada will accept the deal that US proposes.